# PLUM: Improving Code LMs with Execution-Guided On-Policy Preference Learning Driven By Synthetic Test Cases

### v3

@@ 1 Introduction
Language models pre-trained on code corpora have excelled at code generation (Rozière et al., 2024; Li et al., 2023). Supervised Fine-Tuning (SFT) enhances their ability to follow natural language prompts but focuses on reproducing patterns from training data rather than ensuring code correctness (Wei et al., 2023; Zheng et al., 2024). This leads to models that generate syntactically correct but functionally flawed code, unable to meet real-world requirements like edge cases or algorithmic accuracy (Chen et al., 2024).
Works like AlphaCode (Li et al., 2022) and LeTI (Wang et al., 2024b) have introduced test outcomes as a means to define functional correctness in code generation. Building on the insights from these efforts, we propose leveraging preference learning for refining model behavior. Preference learning trains models to prefer certain solutions (e.g., factual, helpful, or harmless) over undesirable ones (e.g., inaccurate, unhelpful, or harmful). Despite its success in aligning models with human values and improving reasoning in other domains (Dong et al., 2023; Guo et al., 2024a; Yuan et al., 2024; Wang et al., 2024a; Pang et al., 2024), the application of preference learning as a principled and efficient approach in code generation remains under-explored, largely due to the lack of high-quality training data.
Recent research shows that reducing the likelihood of incorrect outputs is more effective for improving model performance than simply maximizing correct responses (Setlur et al., 2024; Tajwar et al., 2024). Mode-seeking objectives, which prioritize minimizing errors, have been found to outperform maximum likelihood methods by more efficiently redistributing probability mass across potential outputs. This underscores the importance of applying on-policy and online approaches to enhance preference learning algorithms (Tajwar et al., 2024; Guo et al., 2024b; Setlur et al., 2024; Liu et al., 2024). Unlike offline preference data, on-policy data remains in-distribution with the model, reducing the risk of misalignment (Guo et al., 2024b; Zhang et al., 2024; Tang et al., 2024; Fisch et al., 2024). The main challenge now is how to efficiently obtain preference labels for on-policy data at scale (Yang et al., 2024). In programming tasks, test cases present as a native and powerful candidate solution to address this issue. Being able to automatically produce high-quality test cases unlocks the possibility of collecting preference data over programming questions at any scale.
Prompt for Test Case Generation You are a teaching assistant helping to write reference solutions and tests for programming questions. Given a programming question, you need to first analyze the problem, then write a reference solution (code), followed by assertions that test student solutions. The test code must be runnable when concatenated at the end of student solutions to check the correctness. Programming Question:
@@
Follow the format below: [Analysis]
@@
PLUM utilizes natural language instructions from well-established datasets such as OSS-Instruct (Wei et al., 2023), Evol-Instruct-Code (Luo et al., 2023), and ShareGPT (Cleston, 2023). For each instruction, high-quality test cases are constructed, and multiple solutions are sampled from the model. These solutions are evaluated using the generated test cases, with preference labels assigned based on the results: solutions passing the tests are preferred, while failures are dis-preferred. This dataset then trains the policy using established preference learning algorithms (Rafailov et al., 2023; Ethayarajh et al., 2024; Azar et al., 2023). By relying solely on policy model’s self-generated solutions, PLUM eliminates the need for external, synthetic off-policy data, reducing the risk of distributional shifts and poor generalization commonly observed for synthetic off-policy data, enhancing the model’s robustness and improving its ability to differentiate between correct and incorrect solutions. In addition, by showcasing the effectiveness of our framework, we demonstrated the feasibility of bypassing tedious (and potentially unstable) reward model training (Liu et al., 2023a; Le et al., 2022) and manual labeling, by automating the process of test case synthesis. These simplicity advantages of PLUM makes online preference training of language models possible.
We evaluate PLUM on a diverse set of state-of-the-art code language models under different set-ups, on commonly used evaluation benchmarks: HumanEval(+) and MBPP(+) (Chen et al., 2021; Austin et al., 2021; Liu et al., 2023b) as well as more challenging code generation datasets like LiveCodeBench and LeetCode (Jain et al., 2024; Guo et al., 2024a). We demonstrate that our approach seamlessly integrates with various models in a plug-and-play manner, relying solely on coding instructions to enhance models’ code generation capabilities. Furthermore, we show that online training, facilitated by automated test case generation, further boosts model performance particularly on difficult coding benchmarks, echoing findings from other domains (Xiong et al., 2024b).
@@ 2 Preference Learning Augmented with Test Cases for Code LMs (PLUM)
The core of PLUM lies in leveraging recent advancements in on-policy and online preference learning, which have proven effective across various domains (Xiong et al., 2024b; a; Mitra et al., 2024). In the context of code generation, PLUM simplifies and scales the process of collecting preference data by using test cases. These test cases act as a lightweight yet robust mechanism for evaluating model outputs.
@@
2.1 The PLUM
2.2 Generating Test Cases
Collecting instructions from established datasets
We collect natural language instructions from established datasets including OSS-Instruct (Wei et al., 2023), Evol-Instruct-Code (Luo et al., 2023), and ShareGPT (Cleston, 2023).  [^2] We focus on Python due to its wide use and the availability of well-established training and evaluation resources.These datasets provide a diverse range of programming tasks and instructions. Although they come with gold/silver solutions in the training splits, these solutions are never used in PLUM. Instead, they allow us to directly compare PLUM’s performance against SFT, which relies on gold solutions, as we investigate in our experiments. Not requiring gold solutions for training broadens the applicability of PLUM to a wide variety of real-world coding tasks and user requirements.
@@ 2.2 Generating Test Cases
Given a training instruction in natural language, we prompt a generator model to produce a reference solution, a starter code snippet specifying the function signature, and a suite of test cases using the prompt in Figure 1. The generated test cases are critical for ensuring that the solutions meet the functional requirements specified in the instructions. The correctness of the test cases is central to the success of preference learning. We adopt a consistency-based approach inspired by Chen et al. (2023) and Rozière et al. (2024) for quality control.
@@ 2.3 Sampling Solutions from the Policy to Create the Preference Data
Many preference learning algorithms assume that the preference data is in-distribution for the policy, i.e., the solutions are sampled from the policy model to be trained Rafailov et al. (2023); Ethayarajh et al. (2024); Azar et al. (2023). In practice, however, preference data often contains solutions sampled from different models than the policy, leaving the data out of distribution Bai et al. (2022); Yuan et al. (2024). A common workaround is to first perform supervised fine-tuning (SFT) on the same instructions before applying preference learning (Rafailov et al., 2023; Yuan et al., 2024). This ensures that the policy has a similar distribution to that from which the preference data are sampled.
@@
Moreover, as a recent work points out, training with code snippets containing syntax errors may hurt the model’s performance (Wang et al., 2024b). Solutions passing all test cases are used as the chosen solutions, and those failing at least one the rejected solutions.
@@
This aims to ensure that the learned policy does not drift too far from the original one as drastic changes might cause the model to forget previously learned information or to perform poorly on tasks it was previously adept at Rafailov et al. (2023).
2.4 Preference Learning
We mainly consider two popular preference learning algorithms - Direct Preference Optimization Rafailov et al. (2023) and Kahneman-Tversky Optimization (Ethayarajh et al., 2024) that have been shown to bring improvements for reasoning tasks (Mitra et al., 2024; Yuan et al., 2024; Dubey et al., 2024). For DPO, we subsample redundant classes and randomly pair positive and negative responses for each programming question. In contrast, we use all available responses when training with KTO.
3 Experiments
To demonstrate the effectiveness of PLUM, we evaluate it on established benchmarks: HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and EvalPlus (a widely-adopted augmented version (Liu et al., 2023b) of them). We also use the more challenging LiveCodeBench (Jain et al., 2024).
Datasets
Preference Data Collection
We use GPT-4 (OpenAI et al., 2024) to generate test cases for each programming question. For the OSS-Instruct dataset and ShareGPT dataset, we query GPT-4 for 3 responses for a randomly chosen subset of 1500 questions, and due to the comparatively more complex nature of the natural language instruction, we generate 6 for each of the EvolInstruct instances for a subset of 1000.
Models
We consider a diverse set of strong open language models: MagiCoder (Wei et al., 2023), OpenCodeIntepreter (Zheng et al., 2024), CodeQwen (Bai et al., 2023), DeepSeek Coder (Guo et al., 2024a) and StarCoder2 (Li et al., 2023). MagiCoder and OpenCodeIntepreter contain instruction-tuned checkpoints from DeepSeek Coder and CodeLlama (Rozière et al., 2024) base models. In the main text, we focus on instruction-tuned language models, while differing results from directly training base models to Appendix A.5
Baselines
We evaluate our approach against a variety of baselines, including both prompting-based and fine-tuning techniques. To compare methods that incorporate program correctness through execution feedback, we benchmark our approach against Reflexion (Shinn et al., 2023), using instruction-tuned (SFT) models and value-conditioning techniques (Li et al., 2022; Wang et al., 2024b). Additionally, to contrast preference-learning techniques with the SFT approach, we perform experiments using rejection-sampling-based SFT, utilizing the same set of positive examples as used in KTO. In these experiments, the solutions are also generated on-policy.
3.1 Training
@@ 3.2 Results
Table 3: Comparison with Reflexion (Shinn et al., 2023)
MagiCoder models (-DS, -S-CL, and -S-DS) and OpenCodeIntepreter models (-CL and -DS) have already seen these instructions during supervised fine-tuning, while DeepSeekCoder-Instruct has not, as it was released earlier than the dataset. CodeQwen chat model uses proprietary data. PLUM-ShareGPT data for preference learning is generated with the same setting. Similarly, Table 4 corresponds to the results when we apply PLUM to EvolInstruct (Luo et al., 2023) dataset. Since the instructions are comparatively less clear than the OSS-Instruct dataset, we control the number of initial samples to be the same by generating $50$ samples for each problem and use about 400 instances in total.
@@
We noticed that PLUM-KTO consistently out-performs the baseline, and that PLUM-DPO sometimes under-perform PLUM-KTO. Prior works (Mitra et al., 2024; Yuan et al., 2024) noticed the phenomenon where DPO can exhibit instability due to reducing reward for the positive class.
@@
As shown in Table 3, verbal reinforcement learning like Reflexion (Shinn et al., 2023), does not perform well on code language models fine-tuned for code at the scale relevant to our work. This is partly due to the limitations of smaller LLMs in handling various types of instructions effectively.
Approaches like LeTI (Wang et al., 2024b) implicitly optimizes the model for generating correct programs solely through input prompt, without directly enforcing such distinction with its training objective. Unlike preference learning algorithms ,approaches like LeTI Wang et al. (2024b) optimize models to generate correct programs based solely on the input prompt, without explicitly enforcing correctness through the training objective. As a result, we observe inconsistent outcomes when applying these methods to our tested SFT models, as shown in Table 2. Additionally, our results show that using value-token-conditioned approaches often leads to reduced performance, likely due to the introduction of new value tokens and the absence of clear labels distinguishing good from bad outputs.
@@
We study how important on-policy training is to the performance. The result is presented in 6. To test this, we apply the same method but use preference data sampled from other models (i.e., off-policy). We selected models of varying sizes, including DeepSeek-Coder-1.3B-Instruct, Qwen-2.5-Coder-1.5B-Instruct, CodeStral-22B (MistralAI, 2024), and DeepSeek-Coder-33B-Instruct. We followed the exact same data collection and training processes. Our results show that while off-policy training still improves model performance, it generally underperforms compared to on-policy training. These findings are consistent with prior research  (Xiong et al., 2024a; Dong et al., 2024; Tajwar et al., 2024; Xu et al., 2024).
@@
We experiment with including only non-executable samples as rejected solutions, while using the same set of chosen solutions. As shown in Figure 2, we observe that this is worse than PLUM in most cases. More importantly, it does not always improve the model’s performance and may even hurt. This has also been noted in previous studies (Wang et al., 2024b). Although the positive examples used are the same as our oracle-based preference learning, lower-quality negative examples do not necessarily help the model improve due to the additional noise in the preference signal.
@@
We demonstrate in Table 7 that PLUM enables iterative on-policy alignment while providing accurate preference signals without requiring reward model training. Notably, iterative preference learning, facilitated by the online feedback loops generated through the test case collection procedure, outperforms offline methods on the challenging LeetCode benchmark, using the same data. This finding aligns with results from other domains (Xiong et al., 2024a; b), further confirming the advantages of online learning. This demonstrates the further potential of PLUM in advancing code language models especially in more challenging problems by its support for efficient online policy improvement.
@@ 4 Related Works
Preference learning algorithms like Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Kahneman & Tversky’s Optimization(KTO)  (Ethayarajh et al., 2024) are popular for their cost efficiency and training stability. Beyond controlling model-user interactions, these methods are now applied to more complex reasoning tasks. Ocra-Math (Mitra et al., 2024) uses iterative preference learning to improve math reasoning in SFT-ed language models, while Eurus (Yuan et al., 2024) leverages preference trees for solving complex problems through multi-step interactions with external feedback.
@@
Code generation has become a key application of generative language models. Pre-training on code corpora has led to strong performance in models like StarCoder (Li et al., 2023), StarCoder2 (Lozhkov et al., 2024), and DeepSeek-Coder (Guo et al., 2024a), while others like CodeQwen (Bai et al., 2023) and CodeLlama (Rozière et al., 2024) benefit from continued pre-training on additional code data. To enhance these models, supervised fine-tuning on instruction-response pairs has been employed (Luo et al., 2023; Wei et al., 2023; Zheng et al., 2024). Reinforcement learning techniques, like those in CodeRL (Le et al., 2022; Shojaee et al., 2023; Liu et al., 2023a), and reward models used in DeepSeek-Coder-V2 (DeepSeek-AI et al., 2024) also improve performance using test feedback.
@@
Automated test case generation (Pacheco et al., 2007; Fraser & Arcuri, 2011; Panichella et al., 2015) is crucial for ensuring software quality and safety and has long been a key topic in software engineering. The advent of LLMs has inspired works using transformer models for test generation, either by training models (Tufano et al., 2021; Li et al., 2022) or prompting them (Chen et al., 2023). Test cases also help clarify user intent, aligning model-generated programs with user requirements (Fakhoury et al., 2024; Endres et al., 2024).
@@
Programming-by-examples (Gulwani, 2016) and test-driven programming (Perelman et al., 2014) focus on using test cases to automatically refine programs to meet specifications. This concept has been adapted to enhance deep learning approaches to code synthesis (Kulal et al., 2019; Chen et al., 2023; Zelikman et al., 2023). Recent methods, like CodeT (Chen et al., 2023) and Parsel (Zelikman et al., 2023), use test cases to reduce the search space during inference, while ALGO (Zhang et al., 2023) employs brute-force solutions as oracles to generate test outputs for competitive programming. Our approach, similar to Haluptzok et al. (2023), leverages test cases during training to improve models’ inherent programming capabilities.

### v4

@@ 1 Introduction
Language models pre-trained on code corpora have excelled at code generation [40, 25]. Supervised Fine-Tuning (SFT) enhances their ability to follow natural language prompts but focuses on reproducing patterns from training data rather than ensuring code correctness [49, 58]. This leads to models that generate syntactically correct but functionally flawed code, unable to meet real-world requirements like edge cases or algorithmic accuracy [6].
Works like AlphaCode [26] and LeTI [48] have introduced test outcomes as a means to define functional correctness in code generation. Building on the insights from these efforts, we propose leveraging preference learning for refining model behavior. Preference learning trains models to prefer certain solutions (e.g., factual, helpful, or harmless) over undesirable ones (e.g., inaccurate, unhelpful, or harmful). Despite its success in aligning models with human values and improving reasoning in other domains [10, 19, 54, 47, 36], the application of preference learning as a principled and efficient approach in code generation remains under-explored, largely due to the lack of high-quality training data.
Recent research shows that reducing the likelihood of incorrect outputs is more effective for improving model performance than simply maximizing correct responses [41, 44]. Mode-seeking objectives, which prioritize minimizing errors, have been found to outperform maximum likelihood methods by more efficiently redistributing probability mass across potential outputs. This underscores the importance of applying on-policy and online approaches to enhance preference learning algorithms [44, 20, 41, 29]. Unlike offline preference data, on-policy data remains in-distribution with the model, reducing the risk of misalignment [20, 57, 45, 16]. The main challenge now is how to efficiently obtain preference labels for on-policy data at scale [53]. In programming tasks, test cases present as a native and powerful candidate solution to address this issue. Being able to automatically produce high-quality test cases unlocks the possibility of collecting preference data over programming questions at any scale.
Prompt for Test Case Generation You are a teaching assistant helping to write reference solutions and tests for programming questions. Given a programming question, you need to first analyze the problem, then write a reference solution (code), followed by assertions that test student solutions. The test code must be runnable when concatenated at the end of student solutions to check the correctness.
Programming Question:
@@
Follow the format below:
[Analysis]
@@
PLUM utilizes natural language instructions from well-established datasets such as OSS-Instruct [49], Evol-Instruct-Code [31], and ShareGPT [8]. For each instruction, high-quality test cases are constructed, and multiple solutions are sampled from the model. These solutions are evaluated using the generated test cases, with preference labels assigned based on the results: solutions passing the tests are preferred, while failures are dis-preferred. This dataset then trains the policy using established preference learning algorithms [39, 14, 2]. By relying solely on policy model’s self-generated solutions, PLUM eliminates the need for external, synthetic off-policy data, reducing the risk of distributional shifts and poor generalization commonly observed for synthetic off-policy data, enhancing the model’s robustness and improving its ability to differentiate between correct and incorrect solutions. In addition, by showcasing the effectiveness of our framework, we demonstrated the feasibility of bypassing tedious (and potentially unstable) reward model training [27, 24] and manual labeling, by automating the process of test case synthesis. These simplicity advantages of PLUM makes online preference training of language models possible.
We evaluate PLUM on a diverse set of state-of-the-art code language models under different set-ups, on commonly used evaluation benchmarks: HumanEval(+) and MBPP(+) [7, 1, 28] as well as more challenging code generation datasets like LiveCodeBench and LeetCode [22, 19]. We demonstrate that our approach seamlessly integrates with various models in a plug-and-play manner, relying solely on coding instructions to enhance models’ code generation capabilities. Furthermore, we show that online training, facilitated by automated test case generation, further boosts model performance particularly on difficult coding benchmarks, echoing findings from other domains [51].
@@ 2 Preference Learning Augmented with Test Cases for Code LMs (PLUM)
The core of PLUM lies in leveraging recent advancements in on-policy and online preference learning, which have proven effective across various domains [51, 50, 33]. In the context of code generation, PLUM simplifies and scales the process of collecting preference data by using test cases. These test cases act as a lightweight yet robust mechanism for evaluating model outputs.
@@
2.1 The PLUM
2.2 Generating Test Cases
Collecting instructions from established datasets
We collect natural language instructions from established datasets including OSS-Instruct [49], Evol-Instruct-Code [31], and ShareGPT [8].  [^2] We focus on Python due to its wide use and the availability of well-established training and evaluation resources.These datasets provide a diverse range of programming tasks and instructions. Although they come with gold/silver solutions in the training splits, these solutions are never used in PLUM. Instead, they allow us to directly compare PLUM’s performance against SFT, which relies on gold solutions, as we investigate in our experiments. Not requiring gold solutions for training broadens the applicability of PLUM to a wide variety of real-world coding tasks and user requirements.
@@ 2.2 Generating Test Cases
Given a training instruction in natural language, we prompt a generator model to produce a reference solution, a starter code snippet specifying the function signature, and a suite of test cases using the prompt in Figure 1. The generated test cases are critical for ensuring that the solutions meet the functional requirements specified in the instructions. The correctness of the test cases is central to the success of preference learning. We adopt a consistency-based approach inspired by [5] and [40] for quality control.
@@ 2.3 Sampling Solutions from the Policy to Create the Preference Data
Many preference learning algorithms assume that the preference data is in-distribution for the policy, i.e., the solutions are sampled from the policy model to be trained [39, 14, 2]. In practice, however, preference data often contains solutions sampled from different models than the policy, leaving the data out of distribution [4, 54]. A common workaround is to first perform supervised fine-tuning (SFT) on the same instructions before applying preference learning [39, 54]. This ensures that the policy has a similar distribution to that from which the preference data are sampled.
@@
Moreover, as a recent work points out, training with code snippets containing syntax errors may hurt the model’s performance [48]. Solutions passing all test cases are used as the chosen solutions, and those failing at least one the rejected solutions.
This aims to ensure that the learned policy does not drift too far from the original one as drastic changes might cause the model to forget previously learned information or to perform poorly on tasks it was previously adept at [39].
2.4 Preference Learning
We mainly consider two popular preference learning algorithms - Direct Preference Optimization [39] and Kahneman-Tversky Optimization [14] that have been shown to bring improvements for reasoning tasks [33, 54, 12]. For DPO, we subsample redundant classes and randomly pair positive and negative responses for each programming question. In contrast, we use all available responses when training with KTO.
3 Experiments
To demonstrate the effectiveness of PLUM, we evaluate it on established benchmarks: HumanEval [7], MBPP [1], and EvalPlus (a widely-adopted augmented version [28] of them). We also use the more challenging LiveCodeBench [22].
Datasets
Preference Data Collection
We use GPT-4 [34] to generate test cases for each programming question. For the OSS-Instruct dataset and ShareGPT dataset, we query GPT-4 for 3 responses for a randomly chosen subset of 1500 questions, and due to the comparatively more complex nature of the natural language instruction, we generate 6 for each of the EvolInstruct instances for a subset of 1000.
Models
We consider a diverse set of strong open language models: MagiCoder [49], OpenCodeIntepreter [58], CodeQwen [3], DeepSeek Coder [19] and StarCoder2 [25]. MagiCoder and OpenCodeIntepreter contain instruction-tuned checkpoints from DeepSeek Coder and CodeLlama [40] base models. In the main text, we focus on instruction-tuned language models, while differing results from directly training base models to Appendix A.5
Baselines
We evaluate our approach against a variety of baselines, including both prompting-based and fine-tuning techniques. To compare methods that incorporate program correctness through execution feedback, we benchmark our approach against Reflexion [42], using instruction-tuned (SFT) models and value-conditioning techniques [26, 48]. Additionally, to contrast preference-learning techniques with the SFT approach, we perform experiments using rejection-sampling-based SFT, utilizing the same set of positive examples as used in KTO. In these experiments, the solutions are also generated on-policy.
3.1 Training
@@ 3.2 Results
Table 3: Comparison with Reflexion [42]
MagiCoder models (-DS, -S-CL, and -S-DS) and OpenCodeIntepreter models (-CL and -DS) have already seen these instructions during supervised fine-tuning, while DeepSeekCoder-Instruct has not, as it was released earlier than the dataset. CodeQwen chat model uses proprietary data. PLUM-ShareGPT data for preference learning is generated with the same setting. Similarly, Table 4 corresponds to the results when we apply PLUM to EvolInstruct [31] dataset. Since the instructions are comparatively less clear than the OSS-Instruct dataset, we control the number of initial samples to be the same by generating $50$ samples for each problem and use about 400 instances in total.
@@
We noticed that PLUM-KTO consistently out-performs the baseline, and that PLUM-DPO sometimes under-perform PLUM-KTO. Prior works [33, 54] noticed the phenomenon where DPO can exhibit instability due to reducing reward for the positive class.
@@
As shown in Table 3, verbal reinforcement learning like Reflexion [42], does not perform well on code language models fine-tuned for code at the scale relevant to our work. This is partly due to the limitations of smaller LLMs in handling various types of instructions effectively.
Approaches like LeTI [48] implicitly optimizes the model for generating correct programs solely through input prompt, without directly enforcing such distinction with its training objective. Unlike preference learning algorithms ,approaches like LeTI [48] optimize models to generate correct programs based solely on the input prompt, without explicitly enforcing correctness through the training objective. As a result, we observe inconsistent outcomes when applying these methods to our tested SFT models, as shown in Table 2. Additionally, our results show that using value-token-conditioned approaches often leads to reduced performance, likely due to the introduction of new value tokens and the absence of clear labels distinguishing good from bad outputs.
@@
We study how important on-policy training is to the performance. The result is presented in 6. To test this, we apply the same method but use preference data sampled from other models (i.e., off-policy). We selected models of varying sizes, including DeepSeek-Coder-1.3B-Instruct, Qwen-2.5-Coder-1.5B-Instruct, CodeStral-22B [32], and DeepSeek-Coder-33B-Instruct. We followed the exact same data collection and training processes. Our results show that while off-policy training still improves model performance, it generally underperforms compared to on-policy training. These findings are consistent with prior research  [50, 11, 44, 52].
@@
We experiment with including only non-executable samples as rejected solutions, while using the same set of chosen solutions. As shown in Figure 2, we observe that this is worse than PLUM in most cases. More importantly, it does not always improve the model’s performance and may even hurt. This has also been noted in previous studies [48]. Although the positive examples used are the same as our oracle-based preference learning, lower-quality negative examples do not necessarily help the model improve due to the additional noise in the preference signal.
@@
We demonstrate in Table 7 that PLUM enables iterative on-policy alignment while providing accurate preference signals without requiring reward model training. Notably, iterative preference learning, facilitated by the online feedback loops generated through the test case collection procedure, outperforms offline methods on the challenging LeetCode benchmark, using the same data. This finding aligns with results from other domains [50, 51], further confirming the advantages of online learning. This demonstrates the further potential of PLUM in advancing code language models especially in more challenging problems by its support for efficient online policy improvement.
@@ 4 Related Works
Preference learning algorithms like Direct Preference Optimization (DPO) [39] and Kahneman & Tversky’s Optimization(KTO)  [14] are popular for their cost efficiency and training stability. Beyond controlling model-user interactions, these methods are now applied to more complex reasoning tasks. Ocra-Math [33] uses iterative preference learning to improve math reasoning in SFT-ed language models, while Eurus [54] leverages preference trees for solving complex problems through multi-step interactions with external feedback.
@@
Code generation has become a key application of generative language models. Pre-training on code corpora has led to strong performance in models like StarCoder [25], StarCoder2 [30], and DeepSeek-Coder [19], while others like CodeQwen [3] and CodeLlama [40] benefit from continued pre-training on additional code data. To enhance these models, supervised fine-tuning on instruction-response pairs has been employed [31, 49, 58]. Reinforcement learning techniques, like those in CodeRL [24, 43, 27], and reward models used in DeepSeek-Coder-V2 [9] also improve performance using test feedback.
@@
Automated test case generation [35, 17, 37] is crucial for ensuring software quality and safety and has long been a key topic in software engineering. The advent of LLMs has inspired works using transformer models for test generation, either by training models [46, 26] or prompting them [5]. Test cases also help clarify user intent, aligning model-generated programs with user requirements [15, 13].
@@
Programming-by-examples [18] and test-driven programming [38] focus on using test cases to automatically refine programs to meet specifications. This concept has been adapted to enhance deep learning approaches to code synthesis [23, 5, 55]. Recent methods, like CodeT [5] and Parsel [55], use test cases to reduce the search space during inference, while ALGO [56] employs brute-force solutions as oracles to generate test outputs for competitive programming. Our approach, similar to [21], leverages test cases during training to improve models’ inherent programming capabilities.